dannyward:

The formula for converting degrees Celsius (C) to degrees Fahrenheit (F) is F=9/5C+32 Solve the formula for degrees Celsius (C). Then find the temperature in degrees Celsius (C) when the temperature in degrees Fahrenheit (F) is –4.

Tranquility:

Hello there! Welcome to QuestionCove! Do you have any ideas on how to attempt your question?

Tranquility:

Moving to Mathematics from Arts.

Tranquility:

F = (9/5)C + 32 To solve for C, follow these steps We want to isolate C and get it all by itself. So subtract 32 on both sides of the equation. What do you get? After that, we'll only have 9/5 C on the right hand side. We want to get rid of the 9/5 so that C is all by itself. To do that, we have to multiply by its reciprocal The reciprocal of a/b is b/a When we multiply by the reciprocal, it gets rid of the fraction For example (3/4)a * 4/3 = a Because 3/4 * 4/3 = 1 Following those steps, what do you get when you solve the equation for C?

Tranquility:

@dannyward